OpenStudy (anonymous):

You have a nightlight plugged into an outlet in the hallway, which uses 3.5 watts when plugged in. If the house circuit provides 120.0 volts, what is the current through this bulb?

OpenStudy (anonymous):

Is this a physics question?

OpenStudy (anonymous):

if it is there should b an equation for that

OpenStudy (anonymous):

i think so idk the equation tho

OpenStudy (anonymous):

one sec ill help u

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do uk how to find resistence?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I think what you have to do is divide watts by volts so 3.5/120=0.029166666

OpenStudy (anonymous):

check if that is right

OpenStudy (anonymous):

yeah its right do uk this A circuit is used in which process?Select one of the options below as your answer:A. electricityB. magnetismC. velocityD. gravity

OpenStudy (anonymous):

its electricity

OpenStudy (anonymous):

dang ur good what about thisIn the process of electricity, what flows through the wires?Select one of the options below as your answer:A. protonsB. neutronsC. electronsD. positrons

OpenStudy (anonymous):

electrons?

OpenStudy (anonymous):

can u help me with the rest of my questions i have to answer

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

How much current is in a circuit that includes a 9.0-volt battery and a bulb with a resistance of 4.0 ohms?Select one of the options below as your answer:A. 0.44 ampsB. 36 ampsC. 2.3 ampsD. 13 amps

OpenStudy (anonymous):

Its C.2.3 because the equation for current is I=V/R where I=current, V=voltage, and R=resistance I=9.0/4.0=2.25

OpenStudy (anonymous):

What is the voltage in a circuit that has a current of 10.0 amps and a resistance of 28.5 ohms?Select one of the options below as your answer:A. 2.85 voltsB. 285 voltsC. 0.35 voltsD. 18.5 volts

OpenStudy (anonymous):

Using the equation that i provided you just plug in the number for the corresponding variable and solve for V

OpenStudy (anonymous):

V=I*R V=10.0*28.5 V=285Volts

OpenStudy (anonymous):

i literally just got the answer when u gave it to me haha

OpenStudy (anonymous):

Nice u got it

OpenStudy (anonymous):

What type of relationship exists between the length of a wire and the resistance, if all other factors remain the same? Select one of the options below as your answer:A. Resistance is directly related to length.B. Resistance is directly related to the square of the length.C. Resistance is inversely related to the length.D. Resistance is inversely related to the square of the length.

OpenStudy (anonymous):

I think its A. Resistance is directly related to length. but im not sure

OpenStudy (anonymous):

what does it mean by square of the length?

OpenStudy (anonymous):

idk i answered it n it was correct

OpenStudy (anonymous):

In which situation will the lowest resistance occur?Select one of the options below as your answer:A. thick wire and high temperatureB. thin wire and high temperatureC. thick wire and cold temperatureD. thin wire and low temperature

OpenStudy (anonymous):

Try Thick wire and high Temperature just because it seems logical

OpenStudy (anonymous):

What does the abbreviation DC stand for in physics?Select one of the options below as your answer:A. direct chargeB. dark circuitC. direct currentD. drop charge

OpenStudy (anonymous):

Direct Current

OpenStudy (anonymous):

Which is true regarding AC current?Select one of the options below as your answer:A. The electrons move back and forth.B. The +/- polarity is constant.C. It is common in portable devices.D. It can only produce low voltages.

OpenStudy (anonymous):

I think its A. The electrons move back and forth but i have to double check that

OpenStudy (anonymous):

Think about the electricity sent to your home from a power plant. How does the voltage of the electricity that leaves the plant compare to the voltage of the electricity as it enters your home? Select one of the options below as your answer:A. low voltage to high voltageB. low voltage to low voltageC. high voltage to low voltageD. high voltage to high voltage

OpenStudy (anonymous):

High Voltage to Low Voltage due to resistance

OpenStudy (anonymous):

A battery is the power source for what type of current?Select one of the options below as your answer:A. ACB. BCC. DCD. all types of current

OpenStudy (anonymous):

D. All types of current

OpenStudy (anonymous):

How many question do u have?

OpenStudy (anonymous):

What do these circuits all have in common? Select one of the options below as your answer:A. They all contain switches.B. They are all AC. C. They are all DC.D. They all contain resistors.

OpenStudy (anonymous):

im guessing D

OpenStudy (anonymous):

The symbol below is used to represent what component in a circuit? Select one of the options below as your answer:A. AC power sourceB. DC power sourceC. switchD. resistor

OpenStudy (anonymous):

whatever u sending i cant see the picture

OpenStudy (anonymous):

oh ok now i got it to work, for the previous one the answer should have been A. They all contain Switches

OpenStudy (anonymous):

Which component is used to measure the current in a circuit?Select one of the options below as your answer:A. switchB. resistorC. ammeterD. voltmeter

OpenStudy (anonymous):

w8 dude did u answer the symbol question yet?

OpenStudy (anonymous):

cuz the answer to that one was C.

OpenStudy (anonymous):

yeah i already answered it its all good tho i can miss a few

OpenStudy (anonymous):

k the answer to this one is C. Ammeter

OpenStudy (anonymous):

Which statement is true regarding the circuit schematic below? Select one of the options below as your answer:A. The resistor is connected near the negative terminal of the battery.B. At the present time, the circuit is complete.C. This is a direct current circuit.D. This is an alternating current circuit.

OpenStudy (anonymous):

Idk this one, just guess, A

OpenStudy (anonymous):

What is the equivalent resistance for a parallel circuit that has two resistors: 18.0 ohms and 23.5 ohms?Select one of the options below as your answer:A. 10.2 ohmsB. 20.8 ohmsC. 5.5 ohmsD. 41.5 ohms

OpenStudy (anonymous):

D. 41.5

OpenStudy (anonymous):

A combined circuit has two resistors in parallel (6.0 ohms and 8.0 ohms) and another in series (13.0 ohms). If the power source is 9.0 volts, what will the current be?Select one of the options below as your answer:A. 1.3 ampsB. 0.33 ampsC. 0.55 ampsD. 2.6 amps

OpenStudy (anonymous):

B.0.33

OpenStudy (anonymous):

In a parallel circuit, what affect will adding more resistors to the circuit have?Select one of the options below as your answer:A. The current will increase.B. The current will decrease.C. The current will decrease by a factor proportional to the number of resistors.D. The current is not affected by the number of resistors.

OpenStudy (anonymous):

Current will increase

OpenStudy (anonymous):

alright dude i have to go ur on ur own now

OpenStudy (anonymous):

wait i have only one more question left

OpenStudy (anonymous):

What is the equivalent resistance for a series circuit with three resistors: 19.2 ohms, 24.6 ohms, and 31.2 ohms?Select one of the options below as your answer:A. 36.6 ohmsB. 12.6 ohmsC. 25 ohmsD. 75 ohms